# Large Language Models Hallucination: A Comprehensive Survey

### 09 Oct 2025 v2

@@ 5.3 Embedding-Based Detection
Embedding-based hallucination detection techniques depend on measuring the semantic similarity between input, output, and external references using vector embeddings. These approaches operate under the assumption that a faithful model output should be semantically close to its sources in a shared embedding space. These approaches can be classified into similarity-, gradient-, and spectral-based approaches.
Similarity-based Approaches. Dale et al. [94] used cross-lingual sentence similarity metrics, including LaBSE, LASER, and XNLI. The metrics are used to detect hallucinations in machine translation outputs by comparing the semantic similarity between the source sentence and the translated output. Low similarity scores between the source and translated sentences suggest semantic drift, which indicates possible hallucination. The authors claim that LaBSE predicted hallucinated outputs in the machine translation task better than other sentence embedding models. In another study, contextual embeddings were utilized to detect hallucination in a real-time unsupervised manner [95]. The best results were obtained using the last token’s embedding from the last transformer layer. Nonkes et al. [96] extended semantic embedding approaches by organizing output sentence embeddings as nodes in a similarity graph. Then, a graph attention network was applied to propagate information across semantically similar outputs. Their approach enables robust detection of hallucinations through both the semantic content and the structural relationships in embedding space.
@@
Spectral-Based Approaches. Recent advances extend embedding-based hallucination detection beyond simple similarity. Dasgupta et al. [98] proposed HalluShift, which detects hallucinations by measuring distribution shifts in internal hidden states and attention layers. It achieved substantial gains over prior methods with AUCROC 89.9% on TruthfulQA and 78.6% on HaloScope. In another study, LapEigvals [99] models attention maps as a graph Laplacian, using spectral features. It attained superior results, achieving an AUCROC of 88.9% on the TriviaQA dataset.
@@ 5.4 Learning-Based Detection
Learning-based detection approaches leverage supervised or unsupervised models trained on annotated data to classify LLM outputs as hallucinated or factual. These models are typically trained on annotated datasets or proxy labels and aim to capture patterns indicative of hallucination beyond rule-based or similarity-based techniques. The learning from data feature enables these methods to generalize to diverse hallucination types across domains and tasks.
@@
Supervised-based Methods. ExHalder [100] is an explanation-enhanced hallucination detector for news headlines. This technique adapts NLI knowledge. It combines a reasoning classifier, an explainer module for human-readable justifications, and a hinted classifier that incorporates explanations as additional features. To enable fine-grained detection, Choi et al. [101] introduced RIPA, a token-level hallucination detector trained on synthetic examples generated through knowledge shuffle and partial hallucination strategies. In the partial hallucination approach, only certain words or phrases in a sentence that are factually correct are replaced with false information. This teaches the model to spot hallucinations that appear in specific parts of a sentence rather than across the whole output. RIPA then identifies both the onset; where the first hallucinated token appears, and the span; the full consecutive tokens that make up the hallucinated content. In a related direction, Zhang et al. [102] proposed PRISM, which leverages prompt-guided internal hidden states as features for a supervised hallucination detector. By crafting prompts that enhance the truthfulness structure in the LLM’s internal representations, PRISM achieves good domain generalization, outperforming previous internal-state and token-probability-based baselines in both accuracy and AUROC. Recent research has further addressed scalability and interpretability challenges in hallucination detection. Hu et al. [103] introduced a two-stage framework that combines a small language model for rapid hallucination detection with an LLM-based constrained reasoner that generates detailed explanations for detected cases. Their approach dramatically reduces inconsistencies between detection and explanation, achieving a high F1 score for identifying inconsistent rationales.
Unsupervised and Weakly-based Methods. Lookback Lens [104] is an unsupervised approach that depends on the attention analysis for hallucination detection. By measuring the lookback ratio, which computes how much the model attends to prior context versus its own outputs, this approach identifies contextual hallucinations without the need for labeled data. This method achieves competitive performance compared to supervised detectors like RIPA and supports real-time mitigation during decoding, though it incurs increased inference time and depends on effective sampling.
@@ 5.5 Self-Consistency-based Detection
Self-consistency is an unsupervised technique that improves the reasoning abilities of LLMs by generating several responses to a single prompt and assessing their internal consistency. Self-consistency does not require ground truth references, making it useful in open-ended and low-resource settings. Self-consistency-based techniques can be categorized into answer and question-based methods. Answer-based methods generate multiple answers to the same query using varied decoding hyperparameters, such as temperature and top-k. It then evaluates the model’s consistency across these responses. Question-based methods, on the other hand, generate answers to paraphrased versions of the same question and assess the model’s consistency across these different phrasings. Figure 7 illustrates the difference between answer and question-based self-consistency hallucination detection.
Answer-based Methods. Manakul et al. [106] introduced SelfCheckGPT, which utilizes self-consistency to detect LLMs’ hallucination. This approach examines various techniques, including BERTScore similarity, question-answering consistency, n-gram probabilities, NLI, and prompt-based evaluation to assess the consistency of the generated text. The key idea of this approach is that if a response is factual, then repeated queries to the same prompt with slight randomness should give consistent responses, whereas hallucinated content would give responses with high variability. Li et al. [107] introduced a comprehensive answer evaluation framework called think twice before trusting (T3). This approach prompts the model to generate multiple candidate answers, reflect on each one, and provide justifications. It then compares these answers to assess their relative trustworthiness, aiming to reduce overconfidence in incorrect outputs and improve the reliability of the final prediction. Experiments across multiple tasks and LLMs show that T3 substantially improves hallucination detection AUROC score and calibration compared to prior self-consistency and confidence-based baselines.
@@ 5.6 Detection Challenges
TABLE II: Comparison of hallucination detection techniques. The detection method column indicates whether retrieval-based methods (R.), uncertainty-based (U.), embedding-based (E.), learning-based (L.), or self-consistency (S.) were employed.
Ref Year Task / Domain Detection Method Dataset(s) Metric(s)
R. U. E. L. S.
[81] 2023 Biography generation WikiBio AUC-PR, Spearman & Pearson Corr.
[82] 2024 Information seeking Favabench F1-score
[83] 2024 QA & summarization HaluEval, HotpotQA, CNN/Daily TPR, TNR, Avg Acc.
[86] 2023 Machine translation WMT2018, Newstest2017 AUC-PR
[88] 2023 Summarization WikiBio AUC-PR, Spearman & Pearson Corr.
[90] 2024 Summarization Wikibio, FactCheckGPT, FELM-Science AUROC & AUC-PR
[91] 2025 Summarization WikiBio & NoteSum AUC, Spearman & Pearson Corr.
[87] 2023 QA SciQ AUCROC, Acc.
[92] 2025 QA Private AUC-PR
[93] 2025 QA TriviaQA, SQuAD, NQ, BioASQ AUCROC
[89] 2024 QA & Summarization for TriviaQA, SQuAD, BioASQ, NQ-Open and SVAMP AUROC & AURAC
[94] 2023 Machine translation WMT Hallucination AUC, Precision@R90
[95] 2024 Text generation HELM AUC & Corr.
[98] 2025 QA TruthfulQA, TriviaQA, CoQA, TydiQA AUC, AUC-PR, Acc., Recall
[99] 2025 QA NQopen, TriviaQA, CoQA, SQuADv2, HaluEval, TruthfulQA AUC
[96] 2024 QA, Summarization MSMARCO-QA, FEVER, Wikibio Macro-recall & AUC-PR
[97] 2024 QA, Dialogue, Summarization HADES, HaluEval, wikibio Acc., F1
[100] 2023 News headline generation Private Acc., Precision, Recall, F1
[102] 2024 QA True-False, LogicStruct Acc., AUC
[103] 2024 QA, Summarization FEVER, NHNET, HaluQA, HaluSum Precision, Recall, F1
[104] 2024 QA, Summarization CNN/DM, NQ AUC
[101] 2023 Summarization, Dialogue WoW, SummEval F1, BLEU, RougeL, ChrF, METEOR, Naturalness, Coherence, Groundedness
[105] 2024 QA, Text generation, Code generation WebQA, Ape210K, HumanEval, WordCnt, HaluEval-QA Acc., F1
[106] 2023 Biography generation WikiBio AUC-PR, Factuality Rank
[107] 2024 QA IMDB, CommonsenseQA, PIQA AUC, AUC-PR
[108] 2023 QA TruthfulQA, HaluEval, NQ AUC
[109] 2025 QA TruthfulQA, HotpotQA, FreshQA Precision, Recall, F1
[110] 2025 QA TriviaQA, SQuAD AUC, AUCRAC
@@ 6 Hallucination Mitigation
Hallucination mitigation aims to reduce or prevent the emergence of factually inaccurate, ungrounded, and contextually inconsistent responses in LLMs’ outputs. In contrast to hallucination detection, which concentrates on detecting hallucinatory outputs, mitigation seeks to make modifications to LLMs to deliver accurate responses. Mitigating hallucinations is a crucial priority in the advancement and implementation of LLM to ensure LLM output remains factual, trustworthy, and safe. As shown in Figure 8, this survey categorizes the hallucination mitigation techniques proposed in the literature into prompt-based, retrieval-based, reasoning-based, and model-centric training and adaptation-based techniques.
@@ 6.1 Prompt-Based Techniques
Prompt engineering has emerged as an effective way to guide and regulate the behavior of LLMs. A well-constructed prompt can play an essential role in mitigating hallucinations [117]. The notion of prompt design originated with the advent of GPT-2 [118] and GPT-3 [6], when researchers began exploring the capability of pre-trained LLMs to perform various tasks solely by altering the input prompt, without the need for task-specific fine-tuning. In his section, we classify prompt design strategies for hallucination mitigation into four major categories: template, instruction, tag, and in-context learning (ICL)-based prompts.
@@
Instruction-based prompts. Provide explicit and clear instructions to an LLM to guide its output towards the intended result using natural language. Studies [121, 122] have shown that instruction-based prompting, when coupled with techniques such as CoT reasoning or iterative refinement, significantly reduces hallucinations by constraining model responses to align with facts and rationality. Moreover, LLM fine-tuning on instruction-based datasets enhances its ability to obey complex instructions as well as suppress errors [123]. Kim et al. [124] proposed SELF-EXPERTISE, a method that generates knowledge-based instruction datasets rather than relying on LLMs’ parametric knowledge. It extracts factual knowledge from seed dataset outputs to create structured instruction, input, and output pairs, ensuring factual accuracy. Additionally, system instructions provide explicit guidance to generate logically sound responses, particularly in specialized domains, such as law.
Tag-based prompts. These approaches depend on tailoring LLMs to perform certain tasks by integrating additional information ”tags” into the input. These tags or markers are incorporated into the prompt to direct the model’s focus and response formulation. Feldman et al. [125] conducted a study on the impact of tagged prompting on LLM hallucination mitigation. Their research demonstrated a remarkable success rate of 98.88% in eliminating fabricated information when using context-embedded tags. Similarly, Penkov [126] proposed a method that integrates domain-specific tools, such as BioBERT and ChEBI, with tagged prompting. Their approach focused on anchoring LLM outputs to verifiable facts, particularly in specialized fields, such as biomedicine. The study showed that this combination of techniques could substantially reduce hallucinations by providing a semantic framework to which the model can adhere during response generation.

### 18 Mar 2026 v3

@@ 5.3 Embedding-Based Detection
Embedding-based hallucination detection techniques measure the semantic similarity between input, output, and external references using vector embeddings. These approaches assume that the output of a faithful model should be semantically close to its sources in a shared embedding space. These approaches can be classified into similarity-, gradient-, and spectral-based approaches.
Similarity-based Approaches. Dale et al. [106] used cross-lingual semantic comparison between source and translation to detect hallucinations. They evaluated embedding-based similarity models such as LaBSE and LASER, as well as an XNLI entailment model that measures bidirectional semantic consistency. Low semantic alignment between source and translation indicates semantic detachment, which corresponds to hallucination. Among external methods, LaBSE performed substantially better than LASER. Nonkes et al. [107] extended embedding-based approaches by constructing a semantic similarity graph where each generated sentence is a node and the edges connect semantically close outputs. A Graph Attention Network then performs message passing to learn structural patterns in the embedding space, under the hypothesis that hallucinated generations occupy distinct regions of the latent space.
@@ 5.4 Learning-Based Detection
Learning-based detection approaches leverage trained models to classify LLM outputs as hallucinated or factual. These models are typically trained on annotated datasets or proxy labels and aim to capture patterns indicative of hallucination beyond rule-based or similarity-based techniques. The learning-from-data feature enables these methods to generalize to diverse hallucination types across domains and tasks.
@@
Supervised-based Methods. Choi et al. [109] introduced RIPA, a token-level hallucination detector trained on synthetic examples generated through knowledge shuffle and partial hallucination strategies. In the partial hallucination approach, only certain words or phrases in a sentence that are factually correct are replaced with false information. This teaches the model to spot hallucinations that appear in specific parts of a sentence rather than across the whole output. RIPA then identifies both the onset, where the first hallucinated token appears, and the span, the consecutive set of tokens that make up the hallucinated content. In a related direction, Zhang et al. [110] proposed PRISM, which leverages prompt-guided internal hidden states as features for a supervised hallucination detector. By crafting prompts that enhance the truthfulness structure in the LLM’s internal representations, PRISM achieves good domain generalization, outperforming previous internal-state and token-probability-based baselines in both accuracy and AUROC. In another study, HaluGNN [111] formulates hallucination detection as a graph classification problem using a supervised GNN. Each QA instance is converted into a weighted directed graph where nodes are token hidden states and edges are attention weights, preserving both semantic content and token relationships. A graph neural network is trained on a small labeled set to distinguish factual vs hallucinated responses.
Unsupervised and Weakly-based Methods. Park et al. [112] introduced a truthfulness separator vector (TSV) that is added to internal LLM activations during inference to reshape the embedding space. TSV aims to increase separability between truthful and hallucinated representations without fine-tuning the model. It is trained using a small labeled set and pseudo-labeled unlabeled generations via optimal transport alignment and confidence filtering. In another study, Yamada and Arase [113] detected hallucination by training a lightweight encoder-based model with contrastive (triplet) learning. The learning objective pulls faithful generations closer to the input in embedding space while pushing hallucinated generations farther away. This contrastive objective is combined with a standard classification loss to improve separability between hallucinated and non-hallucinated outputs. At inference time, the model simply compares the input and generated text to predict whether hallucination is present, without retrieval or external knowledge sources. In another study, contextual embeddings were utilized to detect hallucinations in a real-time, unsupervised manner [114]. The authors proposed using the internal hidden states of the LLM during inference. They introduced MIND, a lightweight hallucination classifier trained using pseudo-labeled data automatically constructed from Wikipedia. The method records contextualized token embeddings produced at each decoding step and feeds them into an MLP to detect hallucination. The best results were obtained using the last token’s embedding from the last transformer layer.
@@ 5.5 Self-Consistency-based Detection
Self-consistency is an unsupervised technique that improves LLMs’ reasoning by generating multiple responses to a single prompt and assessing their internal consistency. Self-consistency does not require ground truth references, making it useful in open-ended and low-resource settings. Self-consistency-based techniques can be categorized into answer and question-based methods. Answer-based methods generate multiple answers to the same query using varied decoding hyperparameters, such as temperature and top-k. It then evaluates the model’s consistency across these responses. Question-based methods, on the other hand, generate answers to paraphrased versions of the same question and assess the model’s consistency across these different phrasings. Figure 7 illustrates the difference between answer and question-based self-consistency hallucination detection.
Answer-based Methods. Manakul et al. [116] introduced SelfCheckGPT, which utilizes self-consistency to detect LLMs’ hallucination. This approach examines various techniques, including BERTScore similarity, question-answering consistency, n-gram probabilities, NLI, and prompt-based evaluation to assess the consistency of the generated text. The key idea of this approach is that if a response is factual, repeated queries to the same prompt with slight randomness should yield consistent responses, whereas hallucinated content would yield highly variable responses. Li et al. [117] introduced a comprehensive answer evaluation framework called think twice before trusting (T3). This approach prompts the model to generate multiple candidate answers, reflect on each one, and provide justifications. It then compares these answers to assess their relative trustworthiness, aiming to reduce overconfidence in incorrect outputs and improve the reliability of the final prediction. Experiments across multiple tasks and LLMs show that T3 substantially improves the AUROC for hallucination detection and calibration compared to prior self-consistency and confidence-based baselines. More recently, SelfElicit [118] detected hallucination in long-form text by decomposing them into statements and evaluates each one using calibration-based fact-checking. The model predicts True/False/Not Sure and the confidence of “False” serves as a hallucination score. The approach then elicits reflective thoughts conditioned on this evaluation and stores them in a knowledge hypergraph, which provides contextual semantic information for subsequent statements. A self-consistent, NLI-based conflict-resolution mechanism is applied to detect and mitigate contradictions, preventing hallucination snowballing.
@@ 5.6 Detection Challenges
The observed limitations in each category indicate that no single detection paradigm is sufficient in isolation. Hybrid strategies that combine internal confidence signals, consistency checks, and external evidence grounding are increasingly explored to balance scalability, factual reliability, and robustness across tasks and domains. However, hybrid techniques must also be selected carefully, as combining methods with similar failure modes may not necessarily improve robustness. For example, both uncertainty-based and self-consistency approaches struggle when models produce high-confidence but incorrect outputs. This failure mode is also observed in embedding- and attention-based similarity methods, which rely on internal representations rather than external grounding. The usefulness of the detection technique is highly task- and domain-dependent. For instance, attention- or embedding-based similarity signals tend to be more informative in tasks such as summarization and context-based QA, where hallucinations often manifest as faithfulness errors relative to a given source document. In contrast, for open-domain QA, hallucinations often arise from incorrect world knowledge rather than source misalignment, making internal similarity or attention patterns less reliable indicators of factual correctness. In such cases, retrieval-based or knowledge-grounded verification becomes more critical. These observations suggest that effective hallucination detection requires complementary combinations of methods that address different failure modes, rather than redundant hybrids that amplify shared weaknesses.
@@
TABLE II: Comparison of hallucination detection techniques. Detection categories include retrieval-based (R.), uncertainty-based (U.), embedding-based (E.), learning-based (L.), and self-consistency (S.). Results are reported in AUROC metric; * indicates F1-score, and $\dagger$ indicates Accuracy.
Ref Year Task Detection Result Per Dataset
R. U. E. L. S. WikiBio TruthfulQA TriviaQA SQuAD NQ HaluEval-QA HaluEval-Sum RagTruth CNN/DM HotpotQA
[86] 2023 Bio. gen. 74.2
[87] 2024 Info-seeking
[84] 2025 QA
[89] 2025 QA
[88] 2024 QA, Summ. 72.3* 68.5* 72.1*
[85] 2025 QA, Summ., D2T 83.7* 85.0*
[92] 2023 MT
[94] 2023 Summ. 77.7
[100] 2024 Summ. 90.4*
[101] 2025 Summ. 78.3
[93] 2023 QA
[96] 2025 QA
[97] 2025 QA 90.3 83.7 86.7
[98] 2025 QA 81.1 92.2 81.1
[103] 2025 QA 82.9 88.9 79.5 82.7 87.4
[102] 2024 QA, Summ.
[99] 2024 QA, Summ.
[104] 2026 QA, Summ. 73.3 73.0
[105] 2025 QA, Summ. 96.0 60.1
[95] 2025 QA,Summ., Dlg. 93.0 99.2 95.0 53.0
[106] 2023 MT
[107] 2024 QA, Summ. 63.0
[…]
[110] 2024 QA
[111] 2025 QA 68.8 93.0
[112] 2025 QA 88.7 87.2 78.0
[109] 2023 Summ., Dlg.
[114] 2024 Text gen.
[115] 2024 QA, Text, Code 83.8*
[113] 2025 QA, D2T, Summ. 60.4* 63.6*
[116] 2023 Bio. gen. 80.3
[117] 2024 QA
[119] 2023 QA 77.2 88.0
[120] 2025 QA 71.7
[118] 2025 QA
[121] 2025 QA, Bio. gen.
6 Hallucination Explainability
Hallucination explainability extends hallucination detection by explaining why they occurred and where they came from. It focuses on generating human-understandable reasons for a model’s hallucinations [128]. In this survey, we divide the existing studies into model-internal explainability, which interprets hallucinations through the model’s internal behavior, and evidence-based explainability, which grounds explanations in evidence alignment and reasoning structure.
6.1 Model-Derived Explainability
Model-internal explainability treats hallucinations as a consequence of model-internal dynamics. These methods explain hallucination using signals extracted from model outputs or model-inferred quantities. Uncertainty is currently treated as a component of explainability in LLMs [129]. Huang et al. [130] introduced RePPL, which attributes hallucinations to token-level uncertainty derived from instability in semantic propagation through attention layers and low-confidence generation decisions. In another direction, the hallucination explanation is learned during hallucination detection. HuDEx [131] is trained to justify its hallucination judgments directly using natural language, where it provides reasoning behind these judgments. Similarly, Xie et al. [132] introduced FENCE, which is trained to identify factual errors at the claim level. The model classifies claims as supported, contradictory, or unverified and generates explanatory critiques. Retrieval is used in a supporting role, providing external evidence from search engines, knowledge bases, and knowledge graphs to inform the evaluator’s judgments.
6.2 Grounding Explainability
Grounding explainability refers to methods that explain hallucinations by examining how generated content aligns or fails to align with external evidence. Grounding explainability is usually based on source documents, factual references, or claim decomposition. They identify hallucinations by revealing unsupported statements, contradictions, missing evidence, or broken inference chains, often through evidence highlighting, entailment checking, or structured reasoning verification. One prominent direction for grounding-based hallucination explainability is through a relational mapping of evidence to text. HaluCheck [133] provided explainability through evidence-grounded, sentence-level verification. The proposed technique decomposes model outputs into atomic facts. Then, it retrieves supporting documents from external knowledge sources, and uses NLI models to determine whether each fact is entailed. Hallucinations are highlighted directly, which allows users to see which specific statements lack evidence. Likewise, Chen et al. [134] introduced HaluMap, which constructs a segment-level matrix of entailment and contradiction relations between source inputs and generated text using NLI models. HaluMap produces a heatmap-like visualization that highlights which parts of the output are unsupported or contradicted by the source. To improve robustness, they further propose SelfHaluMap, which is a calibration mechanism that estimates background inconsistency within the source document itself and reduces noise in the final explanation. In another direction, Hu et al. [135] introduced a two-stage framework that combines a small language model for rapid hallucination detection with an LLM-based constrained reasoner that generates detailed explanations for detected cases. Their approach dramatically reduces inconsistencies between detection and explanation, achieving a high F1 score for identifying inconsistent rationales. In another study, Orshansky et al. [136] focused on reasoning-structured explanations. They proposed HalluTree, which decomposes summaries into subclaims and organizes verification results into a hierarchical claim tree. Subclaims are categorized as extractive or inferential. Extractive subclaims are directly verifiable against the source, whereas inferential subclaims requires multi-hop reasoning. Extractive claims are checked with lightweight NLI models. On the other hand, inferential claims initiate a reasoning path, where an LLM proposes a set of supporting facts from the source text, logical or mathematical reasoning, or general knowledge. Later, they are organized into a coherent reasoning chain. An LLM then evaluates the claim’s groundedness using CoT reasoning. Both the supporting facts and the reasoning trace are attached into a verification tree to provide an explicit, explainable rationale. Similarly, Galitsky and Rybalov [137] leveraged information gain with abductive reasoning to explain reasoning-based hallucinations. The proposed technique measures how much a model’s claim shifts the probability distribution away from what is supported by the source. It then checks whether any minimal, plausible hypothesis could logically justify that shift. If no simple abductive explanation exists, or only overly complex ones do, the claim is labeled a hallucination, and the missing or faulty reasoning path is explicitly identified.
6.3 Explainability Challenges
Despite recent progress in developing explainability methods for hallucinations, there remain significant open challenges that limit their effectiveness. Unlike classification labels, there is no universally agreed-upon correct explanation for why a model hallucinates in a given instance. Explanations often rely on proxy signals such as attention patterns, structural irregularities, or alignments with external evidence, which may not fully capture the underlying model’s reasoning. This makes it difficult to evaluate the quality, faithfulness, and completeness of explanations produced by explainability methods. Moreover, in techniques that rely on reasoning trees and topological features, there is a gap between producing explanations that are accurate and faithful to the model’s internal mechanisms and those that are interpretable to humans, especially non-experts. This is because such techniques can be too technical for end users, while simpler explanations may omit key reasoning details, which reduces transparency. Those techniques are also computationally expensive, which limits their practicality for long texts, real-time applications, or large corpora. Furthermore, there is a lack of standardized benchmarks and metrics for evaluating explanation quality. Without agreed standards, comparing different methods and measuring improvement is difficult, which slows progress.
@@ 7 Hallucination Mitigation
Hallucination mitigation aims to reduce or prevent the emergence of factually inaccurate, ungrounded, and contextually inconsistent responses in LLMs’ outputs. In contrast to hallucination detection, which focuses on detecting hallucinatory outputs, mitigation seeks to make modifications to LLMs to deliver accurate responses. Mitigating hallucinations is a crucial priority in the advancement and implementation of LLM to ensure LLM output remains factual, trustworthy, and safe. As shown in Figure 8, this survey categorizes the hallucination mitigation techniques proposed in the literature into prompt-based, retrieval-based, reasoning-based, and model-centric training and adaptation-based techniques. Table III summarizes the mitigation studies with the datasets used in each study and the metrics employed to evaluate their techniques. Moreover, Table V presents results on the four most widely used datasets for hallucination mitigation.
Hallucination mitigation is followed by post‑repair methods, which repair the detected hallucination spans with minimal alteration to the rest of the output. Targeted correction methods operate at the span or entity level. Once specific tokens, entities, or propositions are flagged as hallucinated, the model regenerates or edits only those parts, often under additional constraints or with external evidence [138, 139]. On the other hand, partial regeneration methods relax the granularity from spans to larger segments such as sentences, paragraphs, or specific reasoning steps [138, 140]. The system selectively regenerates only the hallucination segments, while preserving the rest of the answer.
@@ 7.1 Prompt-Based Techniques
Prompt engineering has emerged as an effective way to guide and regulate LLM behavior. A well-constructed prompt can play an essential role in mitigating hallucinations [141]. The notion of prompt design originated with the advent of GPT-2 [142] and GPT-3 [6], when researchers began exploring the capability of pre-trained LLMs to perform various tasks solely by altering the input prompt, without the need for task-specific fine-tuning. In his section, we classify prompt design strategies for hallucination mitigation into four major categories: template, instruction, tag, and in-context learning (ICL)-based prompts.
@@
Instruction-based prompts. Provide explicit and clear instructions to an LLM to guide its output towards the intended result using natural language. Studies [145, 146] have shown that instruction-based prompting, when coupled with techniques such as CoT reasoning or iterative refinement, significantly reduces hallucinations by constraining model responses to align with facts and rationality. Moreover, LLM fine-tuning on instruction-based datasets enhances its ability to follow complex instructions and reduce errors [147]. Kim et al. [148] proposed SELF-EXPERTISE, a method that generates knowledge-based instruction datasets rather than relying on LLMs’ parametric knowledge. It extracts factual knowledge from seed dataset outputs to create structured instruction, input, and output pairs, ensuring factual accuracy. Additionally, system instructions provide explicit guidance for generating logically sound responses, particularly in specialized domains, such as law.
Tag-based prompts. These approaches depend on tailoring LLMs to perform certain tasks by integrating additional information ”tags” into the input. These tags or markers are incorporated into the prompt to direct the model’s focus and response formulation. Feldman et al. [149] conducted a study on the impact of tagged prompting on LLM hallucination mitigation. Their research demonstrated a remarkable success rate of 98.88% in eliminating fabricated information when using context-embedded tags. Similarly, Penkov [150] proposed a method that integrates domain-specific tools, such as BioBERT and ChEBI, with tagged prompting. Their approach focused on anchoring LLM outputs to verifiable facts, particularly in specialized fields, such as biomedicine. The study showed that this combination of techniques could substantially reduce hallucinations by providing a semantic framework for the model to adhere to during response generation.